Opened 9 years ago

Closed 2 years ago

#332 closed flight processing (fixed)

CEH07/04, flight day 102a/2010, Hellifield

Reported by: benj Owned by: emca
Priority: alpha 4 low Milestone: 2010 data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by emca)

Data location: ~arsf/arsf_data/2010/flight_data/uk/CEH07_04-2010_102a_Hellifield

Data arrived from ARSF via network transfer on 12th April 2010.

Scientific objective: Unknown (can't find application)

Priority: a4 - process hyperspectral as a3

PI: Andrew Wilson (& CEH Lancaster)

Collected as a single collection with day 102b (Moorhouse, #333) - information on the same logsheet, combined nav data, etc. Have split into two separate collections for processing.

Sensors:

  • Eagle
  • Hawk
  • Leica LIDAR (08/10/2010)
  • Camera (01/09/2010)

Change History (25)

comment:1 Changed 9 years ago by benj

  • Description modified (diff)

comment:2 Changed 9 years ago by benj

  • Description modified (diff)

Gary says priority is a4 from before they were recorded as high/med/low. Process as a4l for now, since original PI has retired, we've had no other contacts chasing this and that way it won't push our processing schedule back.

comment:3 Changed 9 years ago by benj

  • Description modified (diff)

Andrew has asked that we send him the raw hyperspectral data for this flight and he'll process them. We just need to do the applanix and the Lidar.

Andrew should send us back a copy of the data for archiving when he's done, but in case he doesn't I'm leaving it in the processing queue, it'll just stay at the bottom. If/when Andrew does return the processed data to us it can come off the queue.

comment:4 Changed 9 years ago by benj

  • Description modified (diff)

comment:5 Changed 9 years ago by pmlvis10

IPAS navigation processing done.

comment:6 Changed 9 years ago by ella

Applanix navigation processing has been done. Copied to arsf workspace in Samba.

comment:7 Changed 8 years ago by jb362

  • Description modified (diff)

Started camera processing

comment:8 Changed 8 years ago by jb362

Finished camera processing and waiting for delivery check. Removed photographs 1-15 since they were outside of project area.

comment:9 Changed 8 years ago by ella

Starting delivery check.

comment:10 Changed 8 years ago by ella

Camera delivery looks good apart from:

The file in the eventfile directory doesn't have the letter 'a' after the flight day.

comment:11 Changed 8 years ago by jb362

Camera data delivered 01/09/2010.

comment:12 Changed 8 years ago by jb362

  • Description modified (diff)

comment:13 Changed 8 years ago by emca

Starting LiDAR processing

comment:14 Changed 8 years ago by emca

LiDAR data ready for delivery check at
~/airborne/workspace/CEH07_04-2010_102a_Hellifield/delivery/20100929/CEH07_04

comment:15 Changed 8 years ago by iopa

Delivery checking LiDAR.

comment:16 Changed 8 years ago by iopa

LiDAR delivery check report;
Nothing bad to report, ready to deliver.

Just a note: There is only one logsheet for both 102a and 102b so apart from the 3 flightlines in this delivery it also contains ones belonging to 102b.

comment:17 Changed 8 years ago by emca

  • Description modified (diff)

LiDAR data sent to Andrew Wilson & Clare Rowland 08/10/1010

comment:18 Changed 8 years ago by emca

Camera data sent to C. Rowland 15/10/2010

comment:19 Changed 8 years ago by emca

  • Owner set to emca
  • Status changed from new to assigned

Starting hyperspectral processing

comment:20 Changed 8 years ago by emca

Hyperspectral ready for delivery check at
~airborne/workspace/CEH07_04-2010_102a_Hellifield/delivery/20110405

comment:21 Changed 8 years ago by iopa

Hyperspectral check;

  • lev1 files have incorrect julian day 101 (in filenames, readme including command samples)
  • screenshot files have the same error

Otherwise the data seems ready to deliver.

comment:22 Changed 6 years ago by emca

  • Component changed from Processing: general to Archiving

The hyperspectral data was not archived at NEODC by Andrew Wilson or by ARSF. Claire Rowland @ CEH emailed 11/05/2012 asking about the Eagle and Hawk data.

The data was reprocessed using APL by adbe and made available to Claire via ftp 30/05/2012

Ready for archiving

comment:23 Changed 5 years ago by knpa

Merged and deleted workspace version.

comment:24 Changed 3 years ago by dac

Archiving

Started uploading processed LiDAR and RCD, and raw data to NEODC.

comment:25 Changed 2 years ago by dac

  • Resolution set to fixed
  • Status changed from assigned to closed

Archiving

Data now available from NEODC: http://browse.ceda.ac.uk/browse/neodc/arsf/2010/CEH07_04/CEH07_04-2010_102a_Hellifield

Marking as closed.

Note: See TracTickets for help on using tickets.